the difference of four times a number and seven times the same number is 9. find the number

Answer by Maths68(1474)

Let
Number = x
Four times of x = 4*x = 4x
Seven times of x = 7*x = 7x
Difference of 7x & 4x = 7x-4x which is equal to 9
7x-4x=9
3x=9
Divide by 3 both sides of the equation
3x/3=9/3
x=3
Number is 3
Check
7x-4x=9
Plug in the value of x in above equation
7(3)-4(3) = 9
21-12=9
9=9